OpenStudy (anonymous):

find all solutions of the equation on the interval [0,2pie) cos(x+pie/4)-cos(x-pie/4)=1

OpenStudy (anonymous):

you have to use the sum and differnce formula

OpenStudy (anonymous):

thats right...\[\cos p-\cos q=-2\sin(\frac{p+q}{2})\sin(\frac{p-q}{2})\]

OpenStudy (anonymous):

wrong

OpenStudy (anonymous):

it's right..

OpenStudy (anonymous):

then cos(x+pie/4)-cos(x-pie/4)= -2sin(x)sin(pie/4) = 1 then sin(x)sin(pie/4) = -1/2

OpenStudy (anonymous):

\[\cos (x+a)-\cos (x-a)=-2\sin(\frac{x+a+x-a}{2})\sin(\frac{x+a-x+a}{2})\]\[\cos (x+a)-\cos (x-a)=-2\sin x\sin(a)\]

OpenStudy (anonymous):

can you find answer??

OpenStudy (anonymous):

your a is pi/4

OpenStudy (anonymous):

finally sinx = -(1/root(2))

OpenStudy (anonymous):

so x is (5/4)pi and (7/4)pi

OpenStudy (anonymous):

wow how did you do that

OpenStudy (anonymous):

final eq sinx = -(1/root(2)) means that x is over the angle pie